OpenStudy (anonymous):

6+4(2)*3/10-1=y what is y?

OpenStudy (anonymous):

6 + 4(2)*3/10-1 = 6 + 24/10 - 1 (Remember your order of operations!) Th is is = 6 + 2.4 - 1 = 8.4 - 1 = 7.4